OpenStudy (anonymous):

A spherical object has a diameter of 18 cm. It has a spherical inner core with a diameter of 9 cm. What is the volume of the other layer? Use 3.14 to approximate pi and express your answer in hundredths. ___cm3

OpenStudy (anonymous):

Volume of a sphere = 4/3 pi r^2 First find the volume of the outer sphere. V = 4/3 * pi * (0.09 ^2) = 0.034 m^3 Next find the volume of the inner speher V = 4/3 * pi * (0.045^2) = 0.0085 To find the volume of the other layer, subtract the volume of the core from the volume of the Outer surface. V = 0.034 - 0.0085 = 0.0255 ~ 0.03 m^3

OpenStudy (anonymous):

is that what your looking for?

OpenStudy (anonymous):

Yea and Ive tried it so many times and lose where I am. im not sure what the answer is :/

OpenStudy (amistre64):

V1 - V2 is the volume with a radial amount equal to the differences in the radiuses \[\frac43\pi R^3-\frac43\pi r^3=\frac43\pi [R^3-r^3]\]

OpenStudy (anonymous):

Ive tried that way too but I still dunno what the answer is

OpenStudy (amistre64):

what is the difference in the cubic radiuses? lets start there

OpenStudy (anonymous):

9?

OpenStudy (amistre64):

ok, you might want to reconsider the difference between a diameter and a radius

OpenStudy (amistre64):

18-9 = 9 but that is not quite what we are going for 18 and 9 are diameters, we want radiuses, half the diameter 9 and 4.5 are their radiuses cube them 9^3 and 4.5^3 then subtract them

OpenStudy (anonymous):

-637.875

OpenStudy (amistre64):

good, but you subtracted in a backwards order :) drop the negative sign and we are good

OpenStudy (amistre64):

now multiply that by 4, and divide it by 3

OpenStudy (anonymous):

Sorry my internet went out :/ but I got 850.5

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

Never mind it was 2670.57

OpenStudy (amistre64):

637.875(4) = 2551.5 2551.5/3 = 850.5 the last thing to do is multiply it by 3.14 850.5(3.14) = 2670.57